OpenStudy (anonymous):

A can do a piece of work in 4 hours. B and C can do it in 3 hours. A and C can do it in two hours. How long will B take to do it?

OpenStudy (anonymous):

a can do 1/4 th of the work in 1 hour.......................(1) b and c can do 1/3 of the work in 1 hour..................(2) a and c can do 1/2 of the work in 1 hour....................(3) so, c can do (1/2-1/4)=1/4 of the work in 1 hour alone (from equation 1 and 3).......(4) so, b can do (1/3-1/4) work= 1/12 of the work in 1 hour (from equation 2 and 4) so, b can do it in 12 hours alone

OpenStudy (anonymous):

great thanks

OpenStudy (anonymous):

lol, gimme a medal xD